# Resilient Alerting Protocols for Blockchains

### v1

@@ 5. Maximum Bribery Resistance
Consider an alerting protocol with $n$ nodes where each node stakes $\lambda$}} tokens and the operator funds the contract with c{c} tokens. Consider a node $i$. The maximum amount available in the system comes from all other n1n-1 nodes, which could be slashed, each incurring a penalty of at most $\lambda$}}. In this case, the maximum reward node $i$ can receive by alerting is ${{\lambda}}(n-1)+{c}$.
@@ 6.2. Pure-Strategy Nash Equilibria
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\geq{{\lambda}}(n-1)$, then the action $\mathsf{NoAlert}$ weakly dominates $\mathsf{Alert}$ for node $i$. And $\mathsf{NoAlert}$ is strictly dominant if $\beta_{i}>{{\lambda}}(n-1)$.
@@
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node $i$. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}$.
@@
We showed that if the offered bribe $\beta_{i}\leq{{\lambda}}$, or $\beta_{i}\geq{{\lambda}}(n-1)$, then there exists a dominant strategy for the node $i$. However, we now show that if the bribes offered ${{\lambda}}<\beta_{i}<{{\lambda}}(n-1)$, then no symmetric pure NE exists.
@@ 6.3. Mixed-Strategy Nash Equilibria
From this equation, and since $\tfrac{1}{1+y}\geq 1$ when $y=0$ and $\tfrac{1}{1+y}\geq 0$ otherwise, we get that $\beta_{i}\;\geq\;{{\lambda}}\,(n-1)\,Q_{-i}$. Multiplying both sides by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq{{\lambda}}\,(n-1)\,Q$ for all such $i$. Summing over $i=1,\ldots,n$, we get
@@ 7.1. Sufficient Conditions
Let $P$ be a $\lambda$}}-alerting protocol that runs in constant time and satisfies the three properties. Let $n$ be the number of nodes participating in the protocol. First, the players in the game correspond to the nodes in the protocol and the adversary. Second, each node $i$ can either follow the protocol and alert when needed, or do anything else which is considered as not alerting. By the Alert Capability property, each node can always alert before the alerting period ends. I.e., node $i$’s possible actions are $\mathsf{Alert}$ or $\mathsf{NoAlert}$ as in the simultaneous alerting game. For the adversary, she can offer a bribe $\beta_{i}\geq 0$ to each node $i$. Third, by the Deniability property, no party can distinguish between an execution where a node $i$ alerts and another execution where node $i$ does not alert before the alerting period ends. Therefore, no party can observe other nodes’ actions before the alerting period ends. Thus, each node $i$ chooses its action $a_{i}$ without knowledge of any other node’s choice $a_{j}$. Similarly, the adversary offers bribes without prior knowledge of nodes’ actions. Finally, by the definition of a $\lambda$}}-alerting protocol, if a node deviates and does not alert when needed, it pays a penalty of $\lambda$}}. And by the Reward Distribution property, all nodes that alerted symmetrically share the slashed penalties. Therefore, if at least one node sends a valid alert, all non-alerting nodes pay a penalty of $\lambda$}} and all alerting nodes share the slashed value equally as a reward. For the adversary, if no node alerts, she gains $G$ but pays all bribes. If at least one node alerts, she gains nothing but pays the bribes to non-alerting nodes. Thus, the node’s and adversary’s revenues in the protocol match the players’ utilities defined in Equations 1 and 2, respectively. Since the protocol enforces the strategy space, information structure, and payoffs defined in the game, we conclude it implements the simultaneous alerting game. ∎
@@ 7.2.1. Protocol Description
We require the strict synchrony assumption that writing takes exactly $\Delta_{\sf write}$ steps because the standard synchrony assumption does not specify a model, e.g., probability distribution, over write times less than $\Delta_{\sf write}$. One could consider a simple probabilistic model where a message sent at time $t$ is included in the blockchain at time $t+\Delta_{\sf write}$ with probability $1$, but messages sent at time $t<t^{\prime}<t+\Delta_{\sf write}$ are included at time $t+\Delta_{\sf write}$ with probability $0<p(t)$, decreasing with $t$. It is still feasible to use the Lockstep alerting protocol in such a model, however, we get a tradeoff between the Alert Capability and Deniability properties.
@@ 7.3.3. Protocol Description
Reveal Phase: After the commitment phase concludes, and after $N_{\sf commit}$ blocks were added to the blockchain after the commitment transaction, the node constructs a Proof of Publication $\Pi_{\textit{PoP}}$ consisting of the block headers added to the chain since the last checkpoint that the TEE maintains. It calls $\textsf{Unseal-After}({\gamma},\Pi_{\textit{PoP}})$ locally in the TEE. The TEE verifies the proof and, if sufficiently many blocks have been added to the blockchain, returns the tuple $(m,r)$ containing the message and the nonce and an attestation $A$. The node then sends a reveal transaction containing the action $a$, the proof $\Pi_{\mathsf{Alert}}$, the nonce $r$, and the attestation $A$.
On-Chain Verification: After the reveal phase ends, the smart contract AlertingContract verifies the following for the input from each node $i$: First, that the commitment ${\gamma}_{i}$ was published during the commitment window, i.e., after height $H_{0}$ and before height $H^{\star}$. Then that the attestation $A$ is valid and signed by a trusted TEE (e.g., as in (Chen et al., 2025a; Zhang et al., 2024; socrates1024, 2023; Marlin Protocol Team, 2024)). Finally, that the revealed values $(a,h_{\mathsf{Alert}},r)$ match the commitment ${\gamma}_{i}$, and if $a=\mathsf{Alert}$, the hash of the full proof $\Pi_{\mathsf{Alert}}$ matches $h_{\mathsf{Alert}}$ and that the proof $\Pi_{\mathsf{Alert}}$ itself is valid.
@@ 7.3.4. Implementing the simultaneous alerting game
Since ${\gamma}_{i}^{0}$ and ${\gamma}_{i}^{1}$ are computed using a cryptographic hash function, they cannot help distinguish between executions $E_{0}$ and $E_{1}$. By the integrity and confidentiality properties of the TEE, the internal state (including the action and randomness) is inaccessible to any party, including the host (node $i$) in both executions. By the unforgeability of TEE attestations, in both executions, the node cannot fabricate a signature claiming to be from $T$. By the correctness property of program $T$, the only output path for the randomness and thus the action is the Unseal-After function, which strictly requires a valid PoP for time $t>H^{\star}$. The Unseal-After function takes the commitment handle ${\gamma}_{i}$ and a Proof of Publication $\Pi_{\textit{PoP}}=(C_{B},B_{1},\dots,B_{m})$, a sequence of headers starting from the checkpoint $C_{B}$, where ${\gamma}_{i}$ is included in some block $B_{j}$ and  $j+N_{\sf commit}\geq m$. The TEE verifies that the headers form a valid chain extending from the checkpoint it has, that ${\gamma}_{i}$ is included in a block within this chain, and that at least $N_{\sf commit}$ blocks have been appended after the inclusion block. It returns the secret opening $(m,r)$ only if all these checks pass. By our assumption that no single node nor the adversary can fork the blockchain after a sufficiently recent checkpoint block $C_{B}$, neither node $i$ nor the adversary can create a faster path to revealing the node’s action. Therefore, before $H^{\star}$, no party can obtain the randomness or action from the TEE in either execution. Thus, in both executions, the only observable difference is the commitment ${\gamma}_{i}$, which is indistinguishable between the two executions.
@@ 8.1. Protocol Overview
Instead of enforcing simultaneity, the sequential alerting (illustrated in Figure 3) breaks the alerting period into $n$ equal time intervals called slots, each of length $\Delta_{\sf slot}>\Delta_{\sf write}$, and assigns each node a distinct slot in which it can raise an alert. The protocol assigns slots to nodes according to a permutation $\pi$. Denote by $\pi(s)$ the node in slot $1\leq s\leq n$. And by $\pi^{-1}$ the inverse of the permutation $\pi$, such that $\pi^{-1}(\pi(s))=s$. The function $\pi^{-1}$ maps the node to its assigned slot. In a slight abuse of notation, we write $s_{i}$ instead of $s_{\pi^{-1}(i)}$ for the slot assigned to node $i$.
@@ 8.4. Subgame-Perfect Nash Equilibrium Analysis
We use backward induction over the sequentially ordered set of nodes from slot $n$ to slot $1$ to prove that if the nodes assigned to the first s1s-1 slots did not alert, then the adversary bribes nodes in slots $s,\dots,n$ to not alert if and only if the adversary’s remaining budget $B(s)$ is at least $\sum_{j=s}^{n}{{\lambda}}(j-1)$.
Base Case: Consider the node $i=\pi(n)$ assigned to the last slot $n$. If any previous node in slots $1$ to n1n-1 alerted, then the game would have ended, and node $i$ would not have the opportunity to alert. Otherwise, all previous nodes assigned to slots $1$ to n1n-1 choose not to alert. Thus, by Equation 5, node $i$’s utility is ${{\lambda}}(n-1)$ if it alerts, $a_{i}=\mathsf{Alert}$, or $\beta_{\pi(n)}$ if it does not, $a_{i}=\mathsf{NoAlert}$}.
@@ Appendix B Early Reveal Attacks
In this section, we sketch an example of an attack that an adversary can launch against a simultaneous alerting protocol if nodes can reveal their actions before other nodes have committed to theirs. We show the attack costs $O(n\log n)$. Consider the following adversarial strategy. First, the adversary chooses an arbitrary ordering of the nodes. For simplicity, assume the ordering is by node indices, i.e., node $1$ goes first, then node $2$, and so on. The adversary offers bribes to one node at a time, starting from node $1$.
The adversary offers the first node a bribe of $\beta_{1}={{\lambda}}+\varepsilon>0$, in exchange for choosing $\mathsf{NoAlert}$ and revealing its action immediately. If node $1$ declines, the adversary stops the attack. In which case, all nodes will choose $\mathsf{Alert}$ by Claim 3. And node $1$’s utility will be 0 since no other node is slashed. If, however, node $1$ accepts the bribe, its utility becomes $\varepsilon$, which is strictly better. Therefore, it chooses not to alert, it reveals its action immediately, and the adversary proceeds to node $2$.
The adversary proves to node $2$ that node $1$ chose $\mathsf{NoAlert}$, and offers node $2$ a bribe of $\beta_{2}=\frac{{{\lambda}}}{n-1}+{{\lambda}}+\varepsilon$, in exchange for choosing $\mathsf{NoAlert}$ and revealing its action immediately. If node $2$ declines the bribe, then by Claim 3, all nodes with higher indices will choose $\mathsf{Alert}$, and her utility will be $\frac{{{\lambda}}}{n-1}$. If, however, node $2$ accepts the bribe, her utility is $\beta_{2}=\frac{{{\lambda}}}{n-1}+{{\lambda}}+\varepsilon$ which is stricktly better. Thus, it chooses not to alert, it reveals its action immediately, and the adversary proceeds to node $3$.
@@ C.2. Pure Strategy Nash Equilibria
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\geq{{\lambda}}(n-1)+{c}$, then the action $\mathsf{NoAlert}$ weakly dominates $\mathsf{Alert}$ for node $i$. And $\mathsf{NoAlert}$ is strictly dominant if $\beta_{i}>{{\lambda}}(n-1)+{c}$.
@@
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}+\frac{{c}}{n}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node $i$. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}+\frac{{c}}{n}$.
@@
We showed that if the offered bribe $\beta_{i}\leq{{\lambda}}$, or $\beta_{i}\geq{{\lambda}}(n-1)$, then there exists a dominant strategy for the node $i$. However, we now show that if the bribes offered ${{\lambda}}<\beta_{i}<{{\lambda}}(n-1)$, then no symmetric pure Nash equilibrium exists.
@@ C.3. Mixed Strategy Nash Equilibria
Multiplying Equation 10 by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq({{\lambda}}\,(n-1)+{c})\,Q$ for all such $i$. Summing over $i=1,\ldots,n$, we get

### v2

@@ 5. Maximum Bribery Resistance
Consider an alerting protocol with $n$ nodes where each node stakes $\lambda$}} tokens and the operator funds the contract with c{c} tokens. Consider a node ii. The maximum amount available in the system comes from all other n1n-1 nodes, which could be slashed, each incurring a penalty of at most $\lambda$}}. In this case, the maximum reward node $i$ can receive by alerting is ${{\lambda}}(n-1)+{c}$.
@@ 6.2. Pure-Strategy Nash Equilibria
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\geq{{\lambda}}(n-1)$, then the action $\mathsf{NoAlert}$ weakly dominates $\mathsf{Alert}$ for node ii. And $\mathsf{NoAlert}$ is strictly dominant if $\beta_{i}>{{\lambda}}(n-1)$.
@@
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node ii. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}$.
@@
We showed that if the offered bribe $\beta_{i}\leq{{\lambda}}$, or $\beta_{i}\geq{{\lambda}}(n-1)$, then there exists a dominant strategy for the node ii. However, we now show that if the bribes offered ${{\lambda}}<\beta_{i}<{{\lambda}}(n-1)$, then no symmetric pure NE exists.
@@ 6.3. Mixed-Strategy Nash Equilibria
From this equation, and since $\tfrac{1}{1+y}\geq 1$ when $y=0$ and $\tfrac{1}{1+y}\geq 0$ otherwise, we get that $\beta_{i}\;\geq\;{{\lambda}}\,(n-1)\,Q_{-i}$. Multiplying both sides by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq{{\lambda}}\,(n-1)\,Q$ for all such ii. Summing over $i=1,\ldots,n$, we get
@@ 7.1. Sufficient Conditions
Let $P$ be a $\lambda$}}-alerting protocol that runs in constant time and satisfies the three properties. Let $n$ be the number of nodes participating in the protocol. First, the players in the game correspond to the nodes in the protocol and the adversary. Second, each node $i$ can either follow the protocol and alert when needed, or do anything else which is considered as not alerting. By the Alert Capability property, each node can always alert before the alerting period ends. I.e., node $i$’s possible actions are $\mathsf{Alert}$ or $\mathsf{NoAlert}$ as in the simultaneous alerting game. For the adversary, she can offer a bribe $\beta_{i}\geq 0$ to each node ii. Third, by the Deniability property, no party can distinguish between an execution where a node $i$ alerts and another execution where node $i$ does not alert before the alerting period ends. Therefore, no party can observe other nodes’ actions before the alerting period ends. Thus, each node $i$ chooses its action $a_{i}$ without knowledge of any other node’s choice $a_{j}$. Similarly, the adversary offers bribes without prior knowledge of nodes’ actions. Finally, by the definition of a $\lambda$}}-alerting protocol, if a node deviates and does not alert when needed, it pays a penalty of $\lambda$}}. And by the Reward Distribution property, all nodes that alerted symmetrically share the slashed penalties. Therefore, if at least one node sends a valid alert, all non-alerting nodes pay a penalty of $\lambda$}} and all alerting nodes share the slashed value equally as a reward. For the adversary, if no node alerts, she gains $G$ but pays all bribes. If at least one node alerts, she gains nothing but pays the bribes to non-alerting nodes. Thus, the node’s and adversary’s revenues in the protocol match the players’ utilities defined in Equations 1 and 2, respectively. Since the protocol enforces the strategy space, information structure, and payoffs defined in the game, we conclude it implements the simultaneous alerting game. ∎
@@ 7.2.1. Protocol Description
We require the strict synchrony assumption that writing takes exactly $\Delta_{\sf write}$ steps because the standard synchrony assumption does not specify a model, e.g., probability distribution, over write times less than $\Delta_{\sf write}$. One could consider a simple probabilistic model where a message sent at time $t$ is included in the blockchain at time $t+\Delta_{\sf write}$ with probability 11, but messages sent at time $t<t^{\prime}<t+\Delta_{\sf write}$ are included at time $t+\Delta_{\sf write}$ with probability $0<p(t)$, decreasing with $t$. It is still feasible to use the Lockstep alerting protocol in such a model, however, we get a tradeoff between the Alert Capability and Deniability properties.
@@ 7.3.3. Protocol Description
Reveal Phase: After the commitment phase concludes, and after $N_{\sf commit}$ blocks were added to the blockchain after the commitment transaction, the node constructs a Proof of Publication $\Pi_{\textit{PoP}}$ consisting of the block headers added to the chain since the last checkpoint that the TEE maintains. It calls $\textsf{Unseal-After}({\gamma},\Pi_{\textit{PoP}})$ locally in the TEE. The TEE verifies the proof and, if sufficiently many blocks have been added to the blockchain, returns the tuple $(m,r)$ containing the message and the nonce and an attestation AA. The node then sends a reveal transaction containing the action $a$, the proof $\Pi_{\mathsf{Alert}}$, the nonce $r$, and the attestation AA.
On-Chain Verification: After the reveal phase ends, the smart contract AlertingContract verifies the following for the input from each node $i$: First, that the commitment ${\gamma}_{i}$ was published during the commitment window, i.e., after height $H_{0}$ and before height $H^{\star}$. Then that the attestation AA is valid and signed by a trusted TEE (e.g., as in (Chen et al., 2025a; Zhang et al., 2024; socrates1024, 2023; Marlin Protocol Team, 2024)). Finally, that the revealed values $(a,h_{\mathsf{Alert}},r)$ match the commitment ${\gamma}_{i}$, and if $a=\mathsf{Alert}$, the hash of the full proof $\Pi_{\mathsf{Alert}}$ matches $h_{\mathsf{Alert}}$ and that the proof $\Pi_{\mathsf{Alert}}$ itself is valid.
@@ 7.3.4. Implementing the simultaneous alerting game
Since ${\gamma}_{i}^{0}$ and ${\gamma}_{i}^{1}$ are computed using a cryptographic hash function, they cannot help distinguish between executions $E_{0}$ and $E_{1}$. By the integrity and confidentiality properties of the TEE, the internal state (including the action and randomness) is inaccessible to any party, including the host (node ii) in both executions. By the unforgeability of TEE attestations, in both executions, the node cannot fabricate a signature claiming to be from $T$. By the correctness property of program $T$, the only output path for the randomness and thus the action is the Unseal-After function, which strictly requires a valid PoP for time $t>H^{\star}$. The Unseal-After function takes the commitment handle ${\gamma}_{i}$ and a Proof of Publication $\Pi_{\textit{PoP}}=(C_{B},B_{1},\dots,B_{m})$, a sequence of headers starting from the checkpoint $C_{B}$, where ${\gamma}_{i}$ is included in some block $B_{j}$ and  $j+N_{\sf commit}\geq m$. The TEE verifies that the headers form a valid chain extending from the checkpoint it has, that ${\gamma}_{i}$ is included in a block within this chain, and that at least $N_{\sf commit}$ blocks have been appended after the inclusion block. It returns the secret opening $(m,r)$ only if all these checks pass. By our assumption that no single node nor the adversary can fork the blockchain after a sufficiently recent checkpoint block $C_{B}$, neither node $i$ nor the adversary can create a faster path to revealing the node’s action. Therefore, before $H^{\star}$, no party can obtain the randomness or action from the TEE in either execution. Thus, in both executions, the only observable difference is the commitment ${\gamma}_{i}$, which is indistinguishable between the two executions.
@@ 8.1. Protocol Overview
Instead of enforcing simultaneity, the sequential alerting (illustrated in Figure 3) breaks the alerting period into $n$ equal time intervals called slots, each of length $\Delta_{\sf slot}>\Delta_{\sf write}$, and assigns each node a distinct slot in which it can raise an alert. The protocol assigns slots to nodes according to a permutation $\pi$. Denote by $\pi(s)$ the node in slot $1\leq s\leq n$. And by $\pi^{-1}$ the inverse of the permutation $\pi$, such that $\pi^{-1}(\pi(s))=s$. The function $\pi^{-1}$ maps the node to its assigned slot. In a slight abuse of notation, we write $s_{i}$ instead of $s_{\pi^{-1}(i)}$ for the slot assigned to node ii.
@@ 8.4. Subgame-Perfect Nash Equilibrium Analysis
We use backward induction over the sequentially ordered set of nodes from slot $n$ to slot 11 to prove that if the nodes assigned to the first s1s-1 slots did not alert, then the adversary bribes nodes in slots $s,\dots,n$ to not alert if and only if the adversary’s remaining budget $B(s)$ is at least $\sum_{j=s}^{n}{{\lambda}}(j-1)$.
Base Case: Consider the node $i=\pi(n)$ assigned to the last slot $n$. If any previous node in slots 11 to n1n-1 alerted, then the game would have ended, and node $i$ would not have the opportunity to alert. Otherwise, all previous nodes assigned to slots 11 to n1n-1 choose not to alert. Thus, by Equation 5, node $i$’s utility is ${{\lambda}}(n-1)$ if it alerts, $a_{i}=\mathsf{Alert}$, or $\beta_{\pi(n)}$ if it does not, $a_{i}=\mathsf{NoAlert}$}.
@@ Appendix B Early Reveal Attacks
In this section, we sketch an example of an attack that an adversary can launch against a simultaneous alerting protocol if nodes can reveal their actions before other nodes have committed to theirs. We show the attack costs $O(n\log n)$. Consider the following adversarial strategy. First, the adversary chooses an arbitrary ordering of the nodes. For simplicity, assume the ordering is by node indices, i.e., node 11 goes first, then node 22, and so on. The adversary offers bribes to one node at a time, starting from node 11.
The adversary offers the first node a bribe of $\beta_{1}={{\lambda}}+\varepsilon>0$, in exchange for choosing $\mathsf{NoAlert}$ and revealing its action immediately. If node 11 declines, the adversary stops the attack. In which case, all nodes will choose $\mathsf{Alert}$ by Claim 3. And node 11’s utility will be 0 since no other node is slashed. If, however, node 11 accepts the bribe, its utility becomes $\varepsilon$, which is strictly better. Therefore, it chooses not to alert, it reveals its action immediately, and the adversary proceeds to node 22.
The adversary proves to node 22 that node 11 chose $\mathsf{NoAlert}$, and offers node 22 a bribe of $\beta_{2}=\frac{{{\lambda}}}{n-1}+{{\lambda}}+\varepsilon$, in exchange for choosing $\mathsf{NoAlert}$ and revealing its action immediately. If node 22 declines the bribe, then by Claim 3, all nodes with higher indices will choose $\mathsf{Alert}$, and her utility will be $\frac{{{\lambda}}}{n-1}$. If, however, node 22 accepts the bribe, her utility is $\beta_{2}=\frac{{{\lambda}}}{n-1}+{{\lambda}}+\varepsilon$ which is stricktly better. Thus, it chooses not to alert, it reveals its action immediately, and the adversary proceeds to node 33.
@@ C.2. Pure Strategy Nash Equilibria
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\geq{{\lambda}}(n-1)+{c}$, then the action $\mathsf{NoAlert}$ weakly dominates $\mathsf{Alert}$ for node ii. And $\mathsf{NoAlert}$ is strictly dominant if $\beta_{i}>{{\lambda}}(n-1)+{c}$.
@@
In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}+\frac{{c}}{n}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node ii. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}+\frac{{c}}{n}$.
@@
We showed that if the offered bribe $\beta_{i}\leq{{\lambda}}$, or $\beta_{i}\geq{{\lambda}}(n-1)$, then there exists a dominant strategy for the node ii. However, we now show that if the bribes offered ${{\lambda}}<\beta_{i}<{{\lambda}}(n-1)$, then no symmetric pure Nash equilibrium exists.
@@ C.3. Mixed Strategy Nash Equilibria
Multiplying Equation 10 by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq({{\lambda}}\,(n-1)+{c})\,Q$ for all such ii. Summing over $i=1,\ldots,n$, we get